# Stable Offline Value Function Learning with Bisimulation-based Representations

### v3

@@ Abstract
In reinforcement learning, offline value function learning is the procedure of using an offline dataset to estimate the expected discounted return from each state when taking actions according to a fixed target policy. The stability of this procedure, i.e., whether it converges to its fixed-point, critically depends on the representations of the state-action pairs. Poorly learned representations can make value function learning unstable, or even divergent. Therefore, it is critical to stabilize value function learning by explicitly shaping the state-action representations. Recently, the class of bisimulation-based algorithms have shown promise in shaping representations for control. However, it is still unclear if this class of methods can stabilize value function learning. In this work, we investigate this question and answer it affirmatively. We introduce a bisimulation-based algorithm called kernel representations for offline policy evaluation (krope). krope uses a kernel to shape state-action representations such that state-action pairs that have similar immediate rewards and lead to similar next state-action pairs under the target policy also have similar representations. We show that krope: 1) learns stable representations and 2) leads to lower value error than baselines. Our analysis provides new theoretical insight into the stability properties of bisimulation-based methods and suggests that practitioners can use these methods for stable and accurate evaluation of offline reinforcement learning agents.
@@ 1 Introduction
Learning the value function of a policy is a critical component of many reinforcement learning (rl) algorithms (Sutton & Barto, 2018). While value function learning algorithms such as temporal-difference learning (td) have been successful, they can be unreliable. In particular, the deadly triad, i.e., the combination of off-policy updates, function approximation, and bootstrapping, can make td-based methods diverge (Sutton & Barto, 2018; Tsitsiklis & Van Roy, 1997; Baird, 1995). Function approximation is a critical component of value function learning since it determines the representations of state-action pairs, which in turn defines the space of expressible value functions. Depending on how this value function space is represented, value function learning algorithms may diverge (Ghosh & Bellemare, 2020). That is, the value function learning algorithm may not converge to its fixed-point, or may even diverge away from it. In this work, we investigate how to explicitly learn state-action representations to stabilize value function learning.
@@
In seeking such representations, we turn to $\pi$-bisimulation algorithms. These algorithms define a metric to capture behavioral similarity between state-action pairs such that similarity is based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$ (Castro, 2019). The algorithms then use this metric to learn representations such that state-action pairs that are similar under this metric have similar representations. Ultimately, the goal of $\pi$-bisimulation methods is to learn representations such that state-actions pairs with similar values under $\pi$ also have similar representations (see Figure 1). While these algorithms have shown promise in improving the expected return of rl algorithms, it remains unclear whether they contribute to stability (Castro et al., 2023; Zhang et al., 2021; Castro et al., 2022). In this paper, we aim to understand whether the $\pi$-bisimulation-based representations stabilize value function learning.
@@
    We empirically validate that krope representations lead to more stable and accurate offline value function learning compared to baselines (Section 4).
@@
    We empirically analyze the sensitivity of the krope learning procedure under the deadly triad. These experiments shed light on when representation pre-training may be easier than direct value function learning with lspe (Appendix C.3.1).
@@ 2.3 Stable Representations and $q^{\pi_{e}}$-Consistency
Proposition 1 (Asadi et al. 2024; Wang et al. 2021a).
@@ 2.4 Related Works
There are several works that have shown shaping representations can be effective for offline rl (Yang & Nachum, 2021; Islam et al., 2023; Nachum & Yang, 2024; Zang et al., 2023a; Arora et al., 2020; Uehara et al., 2021; Chen & Jiang, 2019; Pavse & Hanna, 2023b). Ghosh & Bellemare (2020) presented a theoretical understanding of how various representations can stabilize td learning. However, they did not discuss bisimulation-based representations. Kumar et al. (2021); Ma et al. (2024); He et al. (2024) promote the stability of td-based methods by increasing the rank of the representations to prevent representation collapse. However, as we show in Section 4, these types of representations can still lead to inaccurate ope. On the other hand, krope mitigates representation collapse and leads to accurate ope. Chang et al. (2022) introduced bcrl to learn Bellman complete representations for stable ope. While in theory, bc representations are desirable, we found that bcrl is sensitive to hyperparameter tuning. In contrast, we show that krope is more robust to hyperparameter tuning. Pavse & Hanna (2023a) showed that bisimulation-based representations mitigate the divergence of fqe; however, they did not provide an explanation for divergence mitigation. Our work provides theoretical insight into the stability properties of bisimulation-based algorithms.
@@
Recently, there has been lot of interest in $\pi$-bisimulation algorithms for better generalization (Ferns et al., 2004, 2011; Ferns & Precup, 2014; Castro, 2019; Zang et al., 2023b). These algorithms measure similarity between two state-action pairs based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$. These algorithms first define a distance metric that captures this $\pi$-bisimilarity, and then use this metric to learn representations such that $\pi$-bisimular states have similar representations (Castro et al., 2022; Castro, 2019; Zhang et al., 2021; Castro et al., 2023; Chen & Pan, 2022; Kemertas & Jepson, 2022). Castro (2019) introduced a $\pi$-bisimulation learning algorithm but assumed that the transition dynamics are deterministic. Zhang et al. (2021) introduced a bisimulation algorithm that allowed for Gaussian dynamics and demonstrated its effectiveness in discarding distracting features for control. Castro et al. (2022, 2023) introduced bisimulation-based algorithms for control that allow for stochastic transition dynamics. To the best of our knowledge, there is still a gap in the literature as to whether the general class of these learned $\pi$-bisimulation-based representations stabilize offline value function learning. To address this gap, we introduce krope, an algorithm that leverages Castro et al. (2023)’s kernel perspective on similarity metrics. We first explicitly define a kernel that captures the relationship between state-action features in latent space in terms of the $\pi$-bisimulation relation. This definition allows us to immediately establish stability since we can now analyze the spectral properties of offline value function learning algorithms. The proofs for krope’s basic theoretical properties (Section 3.1) follow those by Castro et al. (2023). Our stability-related theoretical results (Sections 3.2 and 3.3) and empirical analysis (Section 4) are novel to this work.
@@ 3 Kernel Representations for Offline Policy Evaluation
Prior works typically define a distance metric to capture the notion of $\pi$-bisimilarity between two states: the distance between states is in terms of the immediate rewards received and differences of next states under $\pi$ (see Figure 1) (Castro, 2019). In this work, we follow Castro et al. (2023) and take a perspective of kernels instead of distances. We first define a kernel that captures $\pi$-bisimilarity $k^{{\pi_{e}}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, but for pairs of state-actions and under ${\pi_{e}}$:
@@
By definition $k_{2}$ is a function of independently coupling $(s_{1}^{\prime},a_{1}^{\prime})$ and $(s_{2}^{\prime},a_{2}^{\prime})$ (Castro et al., 2022, 2023). The advantage of this independence is that it enables a practical algorithm and is flexible as it make no assumptions on the nature of the transition dynamics of $P^{\pi_{e}}$. While independent coupling does raise a complication with stochastic dynamics and/or stochastic policies since self-similarity may be lower than similarity between different state-actions, i.e., $k^{\pi_{e}}(s_{1},a_{1};s_{1},a_{1})<k^{\pi_{e}}(s_{1},a_{1};s_{2},a_{2})$ for some $(s_{1},a_{1}),(s_{2},a_{2})\in\mathcal{X}$, $k^{\pi_{e}}$ still induces a valid $\pi$-bisimulation-based metric. To see its validity, consider the induced distance function by $k^{\pi_{e}}$, i.e., $\forall x,y\in\mathcal{X},d_{\text{{krope}}}(x,y):=k^{{\pi_{e}}}(x,x)+k^{{\pi_{e}}}(y,y)-2k^{{\pi_{e}}}(x,y)$. The metric $d_{\text{{krope}}}$ satisfies continuity (see Lemma 3 in Appendix B.1), which is an important property for metrics to satisfy (Lan et al., 2021). Lemma 3 states that the absolute action-value difference between any two state-action pairs under ${\pi_{e}}$ is upper-bounded by $d_{\text{{krope}}}$ plus an additive constant, where the additive constant arises due to the independent coupling. Thus far, prior work has removed the additive constant by assuming deterministic or Gaussian dynamics (Castro, 2019; Zhang et al., 2021). However, these assumptions are practically more restrictive. While one might expect the existence of the additive constant to hurt performance, prior work and our work (see Section 4) show that independent coupling actually yields strong performance (Pavse & Hanna, 2023a; Castro et al., 2023). Furthermore, in Section 3.2 we show that this property does not affect the stability properties of the learned representations. An interesting future direction will be to develop practically feasible approaches to eliminate the additive constant.
@@ 3.3 Connection to Bellman Completeness
In this section, we draw a novel connection between krope representations and Bellman completeness. Chen & Jiang (2019) proved a similar result but focused on bisimulation representations instead of krope representations, which are $\pi$-bisimulation-like representations. We say a function class $\mathcal{F}$ is Bellman complete if it is complete under the Bellman operator: $\mathcal{T}^{{\pi_{e}}}f\subseteq\mathcal{F},\forall f\in\mathcal{F}$. For instance, suppose $\mathcal{F}$ is the class of linear functions spanned by $\Phi$, :=$\mathcal{F}:=\{f\in\mathbb{R}^{\mathcal{X}}:f:=\Phi w\},w\in\mathbb{R}^{d}$. Then if $\mathcal{T}^{{\pi_{e}}}f,\forall f\in\mathcal{F}$ is also a linear function within the span of $\Phi$, we say $\Phi$ is a Bellman complete representation. Bellman completeness is an alternative condition for stability and is typically assumed to ensure to data-efficient policy evaluation (Wang et al., 2021b; Szepesvári & Munos, 2005; Chang et al., 2022).
@@
We now present our second main result. It states that krope representations are Bellman complete:
@@
Takeaway #$\#1$: Stability of Bisimulation-based Representations krope representations induce non-expansive value function updates and are Bellman complete. They avoid divergence of offline value function learning.
@@
where $\mathcal{K}_{\bar{\omega}}(s_{1},a_{1};s_{2};a_{2})=1-\frac{\left|r(s_{1},a_{1})-r(s_{2},a_{2})\right|}{\left|r_{\text{max}}-r_{\text{min}}\right|}+\gamma\mathbb{E}_{{\pi_{e}}}[\tilde{k}_{\bar{\omega}}(s_{1}^{\prime},a_{1}^{\prime};s_{2}^{\prime},a_{2}^{\prime})]$, the state-action pairs are sampled from $\mathcal{D}$, and $\bar{\omega}$ are weights of the target network that are periodically copied from $\omega$ (Mnih et al., 2015). In this work, we use krope as an auxiliary task, which introduces only an auxiliary task weight as the additional hyperparameter.
@@ 4.1 Empirical Setup
Our primary representation learning baseline is fitted q-evaluation (fqe) (Le et al., 2019). fqe is the most fundamental deep rl ope algorithm that learns representations of state-actions to predict the long-term performance of a policy. While fqe is typically used as an ope algorithm, it can also be viewed as a value-predictive representation learning algorithm (Lehnert & Littman, 2020). More specifically, consider its loss function: $\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}}\left[\left(r(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim{\pi_{e}}}[q_{\bar{\xi}}(s^{\prime},a^{\prime})]-q_{\xi}(s,a)\right)^{2}\right]$, where $q_{\xi}(s,a):=\phi_{\xi^{\prime}}(s,a)^{\top}w$ and $\xi=\{\xi^{\prime},w\}$. We view $\xi$ as the neural network weights of an action-value neural network and $w$ as the linear weights of the network applied on the output of the penultimate layer $\phi_{\xi^{\prime}}(s,a)$ of the neural network. Then minimizing this loss function shapes the representations $\phi_{\xi^{\prime}}(s,a)$ to predict the expected future discounted return. As noted in Section 2, we follow the linear evaluation protocol where $\phi_{\xi^{\prime}}$ is shaped by different auxiliary tasks and is then used with lspe for ope since it helps us understand the properties of the representations within the context of a well-understood value function learning algorithm (Grill et al., 2020; Chang et al., 2022; Farebrother et al., 2024; Wang et al., 2021a). We provide the pseudocode of this setup in Appendix A. We also note that in Appendix C, we present results of performing ope using fqe instead lspe, and find that krope still reliably produces stable ope estimates.
We consider the following three classes (and total $6$) auxiliary representation learning algorithms that are typically paired with fqe for stability: I) bisimulation-metric based, II) model-based, and III) co-adaptation penalty-based. In class I, we consider 1) krope (ours), 2) deep bisimulation for control (dbc) (Zhang et al., 2021), 3) representations for ope (rope) (Pavse & Hanna, 2023a; Castro et al., 2022); in class II, we consider bellman complete representation learning (bcrl-exp-na) (Chang et al., 2022); and in class III, we consider 1) absolute dr3 regularizer (Kumar et al., 2021; Ma et al., 2024) and 2) beer regularizer (He et al., 2024). In all cases, the penultimate layer features of fqe’s action-value encoder $\phi_{\xi^{\prime}}$ are fed into lspe for ope. Note that since bcrl was not designed as an auxiliary task (Chang et al., 2022), we evaluate it as a non-auxiliary (na) task algorithm. We provide additional details on the baselines in Appendix C.

### v4

@@ Abstract
In reinforcement learning, offline value function learning is the procedure of using an offline dataset to estimate the expected discounted return from each state when taking actions according to a fixed target policy. The stability of this procedure, i.e., whether it converges to its fixed-point, critically depends on the representations of the state-action pairs. Poorly learned representations can make value function learning unstable, or even divergent. Therefore, it is critical to stabilize value function learning by explicitly shaping the state-action representations. Recently, the class of bisimulation-based algorithms have shown promise in shaping representations for control. However, it is still unclear if this class of methods can stabilize value function learning. In this work, we investigate this question and answer it affirmatively. We introduce a bisimulation-based algorithm called kernel representations for offline policy evaluation (krope). krope uses a kernel to shape state-action representations such that state-action pairs that have similar immediate rewards and lead to similar next state-action pairs under the target policy also have similar representations. We show that krope: 1) learns stable representations and 2) leads to lower value error than baselines. Our analysis provides new theoretical insight into the stability properties of bisimulation-based methods and suggests that practitioners can use these methods to improve the stability and accuracy of offline evaluation of reinforcement learning agents.
@@ 1 Introduction
Learning the value function of a policy is a critical component of many reinforcement learning (rl) algorithms (Sutton & Barto, 2018). While value function learning algorithms such as temporal-difference learning (td) have been successful, they can be unreliable. In particular, the deadly triad, i.e., the combination of off-policy updates, function approximation, and bootstrapping, can make td-based methods diverge (Sutton & Barto, 2018; Tsitsiklis & Van Roy, 1997; Baird, 1995; Hasselt et al., 2018). Function approximation is a critical component of value function learning since it determines the representations of state-action pairs, which in turn defines the space of expressible value functions. Depending on how this value function space is represented, value function learning algorithms may diverge (Ghosh & Bellemare, 2020). That is, the value function learning algorithm may not converge to its fixed-point, or may even diverge away from it. In this work, we explicitly learn state-action representations to improve the stability of offline value function learning.
@@
In seeking such representations, we turn to $\pi$-bisimulation algorithms. These algorithms define a metric to capture behavioral similarity between state-action pairs such that similarity is based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$ (Castro, 2020). The algorithms then use this metric to learn representations such that state-action pairs that are similar under this metric have similar representations. Ultimately, the goal of $\pi$-bisimulation methods is to learn representations such that state-actions pairs with similar values under $\pi$ also have similar representations (see Figure 1). While these algorithms have shown promise in improving the expected return of rl algorithms, it remains unclear whether they contribute to stability (Castro et al., 2023; Zhang et al., 2021; Castro et al., 2021). In this paper, we aim to understand whether the $\pi$-bisimulation-based representations stabilize value function learning.
@@
    We empirically validate that krope representations improve the stability and accuracy of offline value function learning algorithms on $10/13$ offline datasets and against $7$ baselines (Section 4).
@@
    We empirically analyze the sensitivity of the krope learning procedure under the deadly triad. These experiments shed light on when representation learning may be easier than value function learning (Section 4.4).
@@ 2.3 Stable Representations and $q^{\pi_{e}}$-Consistency
Proposition 1 (Asadi et al. 2024; Wang et al. 2021).
@@ 2.4 Related Works
There are several works that have shown shaping representations can be effective for offline rl (Yang & Nachum, 2021; Islam et al., 2023; Nachum & Yang, 2024; Zang et al., 2023a; Arora et al., 2020; Uehara et al., 2022; Chen & Jiang, 2019; Pavse & Hanna, 2023b). Ghosh & Bellemare (2020) presented a theoretical understanding of how various representations can stabilize td learning. However, they did not discuss bisimulation-based representations. Kumar et al. (2022); Ma et al. (2024); He et al. (2024) promote the stability of td-based methods by increasing the rank of the representations to prevent representation collapse. However, as we show in Section 4, these types of representations can still lead to inaccurate ope. On the other hand, krope mitigates representation collapse and leads to accurate ope. Chang et al. (2022) introduced bcrl to learn Bellman complete representations for stable ope. While in theory, bc representations are desirable, we found that bcrl is sensitive to hyperparameter tuning. In contrast, we show that krope is more robust to hyperparameter tuning. Pavse & Hanna (2023a) showed that bisimulation-based representations mitigate the divergence of fqe; however, they did not provide an explanation for divergence mitigation. Our work provides theoretical insight into the stability properties of bisimulation-based algorithms.
@@
Recently, there has been lot of interest in $\pi$-bisimulation algorithms for better generalization (Ferns et al., 2004, 2011; Ferns & Precup, 2014; Castro, 2020; Zang et al., 2023b). These algorithms measure similarity between two state-action pairs based on immediate rewards received and the similarity of next state-action pairs visited by $\pi$. These algorithms first define a distance metric that captures this $\pi$-bisimilarity, and then use this metric to learn representations such that $\pi$-bisimular states have similar representations (Castro et al., 2021; Castro, 2020; Zhang et al., 2021; Castro et al., 2023; Chen & Pan, 2022; Kemertas & Jepson, 2022). Castro (2020) introduced a $\pi$-bisimulation learning algorithm but assumed that the transition dynamics are deterministic. Zhang et al. (2021) introduced a bisimulation algorithm that allowed for Gaussian dynamics and demonstrated its effectiveness in discarding distracting features for control. Castro et al. (2021, 2023) introduced bisimulation-based algorithms for control that allow for stochastic transition dynamics. To the best of our knowledge, there is still a gap in the literature as to whether the general class of these learned $\pi$-bisimulation-based representations stabilize offline value function learning. To address this gap, we introduce krope, an algorithm that leverages Castro et al. (2023)’s kernel perspective on similarity metrics. We first explicitly define a kernel that captures the relationship between state-action features in latent space in terms of the $\pi$-bisimulation relation. This definition allows us to immediately establish stability since we can now analyze the spectral properties of offline value function learning algorithms. The proofs for krope’s basic theoretical properties (Section 3.1) follow those by Castro et al. (2023). Our stability-related theoretical results (Sections 3.2 and 3.3) and empirical analysis (Section 4) are novel to this work.
@@ 3 Kernel Representations for Offline Policy Evaluation
Prior works typically define a distance metric to capture the notion of $\pi$-bisimilarity between two states: the distance between states is in terms of the immediate rewards received and differences of next states under $\pi$ (see Figure 1) (Castro, 2020). In this work, we follow Castro et al. (2023) and take a perspective of kernels instead of distances. We first define a kernel that captures $\pi$-bisimilarity $k^{{\pi_{e}}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, but for pairs of state-actions and under ${\pi_{e}}$:
@@
By definition $k_{2}$ is a function of independently coupling $(s_{1}^{\prime},a_{1}^{\prime})$ and $(s_{2}^{\prime},a_{2}^{\prime})$ (Castro et al., 2021, 2023). The advantage of this independence is that it enables a practical algorithm and is flexible as it make no assumptions on the nature of the transition dynamics of $P^{\pi_{e}}$. While independent coupling does raise a complication with stochastic dynamics and/or stochastic policies since self-similarity may be lower than similarity between different state-actions, i.e., $k^{\pi_{e}}(s_{1},a_{1};s_{1},a_{1})<k^{\pi_{e}}(s_{1},a_{1};s_{2},a_{2})$ for some $(s_{1},a_{1}),(s_{2},a_{2})\in\mathcal{X}$, $k^{\pi_{e}}$ still induces a valid $\pi$-bisimulation-based metric. To see its validity, consider the induced distance function by $k^{\pi_{e}}$, i.e., $\forall x,y\in\mathcal{X},d_{\text{{krope}}}(x,y):=k^{{\pi_{e}}}(x,x)+k^{{\pi_{e}}}(y,y)-2k^{{\pi_{e}}}(x,y)$. The metric $d_{\text{{krope}}}$ satisfies continuity (see Lemma 3 in Appendix B.1), which is an important property for metrics to satisfy (Lan et al., 2021). Lemma 3 states that the absolute action-value difference between any two state-action pairs under ${\pi_{e}}$ is upper-bounded by $d_{\text{{krope}}}$ plus an additive constant, where the additive constant arises due to the independent coupling. Thus far, prior work has removed the additive constant by assuming deterministic or Gaussian dynamics (Castro, 2020; Zhang et al., 2021). However, these assumptions are practically more restrictive. While one might expect the existence of the additive constant to hurt performance, prior work and our work (see Section 4) show that independent coupling actually yields strong performance (Pavse & Hanna, 2023a; Castro et al., 2023). Furthermore, in Section 3.2 we show that this property does not affect the stability properties of the learned representations. An interesting future direction will be to develop practically feasible approaches to eliminate the additive constant.
@@ 3.3 Connection to Bellman Completeness
In this section, we draw a novel connection between krope representations and Bellman completeness. Chen & Jiang (2019) proved a similar result but focused on bisimulation representations instead of krope representations, which are $\pi$-bisimulation-like representations. We say a function class $\mathcal{F}$ is Bellman complete if it is complete under the Bellman operator: $\mathcal{T}^{{\pi_{e}}}f\subseteq\mathcal{F},\forall f\in\mathcal{F}$. For instance, suppose $\mathcal{F}$ is the class of linear functions spanned by $\Phi$, :=$\mathcal{F}:=\{f\in\mathbb{R}^{\mathcal{X}}:f:=\Phi w\},w\in\mathbb{R}^{d}$. Then if $\mathcal{T}^{{\pi_{e}}}f,\forall f\in\mathcal{F}$ is also a linear function within the span of $\Phi$, we say $\Phi$ is a Bellman complete representation. Bellman completeness is an alternative condition for stability and is typically assumed to ensure to data-efficient policy evaluation (Wang et al., 2021; Szepesvári & Munos, 2005; Chang et al., 2022).
@@
We now present our second main result. It states that krope representations are Bellman complete under a reward function injectivity assumption (see proof in Appendix B.2):
@@
Takeaway #$\#1$: Stability of Bisimulation-based Representations Under the theoretical assumptions made, krope representations avoid divergence of offline value function learning. They 1) induce non-expansive value function updates and 2) are Bellman complete.
@@
where $\mathcal{K}_{\bar{\omega}}(s_{1},a_{1};s_{2};a_{2})=1-\frac{\left|r(s_{1},a_{1})-r(s_{2},a_{2})\right|}{\left|r_{\text{max}}-r_{\text{min}}\right|}+\gamma\mathbb{E}_{{\pi_{e}}}[\tilde{k}_{\bar{\omega}}(s_{1}^{\prime},a_{1}^{\prime};s_{2}^{\prime},a_{2}^{\prime})]$, the state-action pairs are sampled from $\mathcal{D}$, and $\bar{\omega}$ are weights of the target network that are periodically copied from $\omega$ (Mnih et al., 2015). In this work, we use krope as an auxiliary task, which introduces only an auxiliary task weight as the additional hyperparameter. It is critical to note that since the learning algorithm is a semi-gradient method, it may still diverge. Nevertheless, in Section 4 we show krope still improves the stability and accuracy of ope compared to baselines.
@@ 4.1 Empirical Setup
Our primary representation learning baseline is fitted q-evaluation (fqe) (Le et al., 2019). fqe is the most fundamental deep rl ope algorithm that learns representations of state-actions to predict the long-term performance of a policy. While fqe is typically used as an ope algorithm, it can also be viewed as a value-predictive representation learning algorithm (Lehnert & Littman, 2020). More specifically, consider its loss function: $\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}}\left[\left(r(s,a)+\gamma\mathbb{E}_{a^{\prime}\sim{\pi_{e}}}[q_{\bar{\xi}}(s^{\prime},a^{\prime})]-q_{\xi}(s,a)\right)^{2}\right]$, where $q_{\xi}(s,a):=\phi_{\xi^{\prime}}(s,a)^{\top}w$ and $\xi=\{\xi^{\prime},w\}$. We view $\xi$ as the neural network weights of an action-value neural network and $w$ as the linear weights of the network applied on the output of the penultimate layer $\phi_{\xi^{\prime}}(s,a)$ of the neural network. Then minimizing this loss function shapes the representations $\phi_{\xi^{\prime}}(s,a)$ to predict the expected future discounted return. As noted in Section 2, we follow the linear evaluation protocol where $\phi_{\xi^{\prime}}$ is shaped by different auxiliary tasks and is then used with lspe for ope since it helps us understand the properties of the representations within the context of a well-understood value function learning algorithm (Grill et al., 2020; Chang et al., 2022; Farebrother et al., 2024; Wang et al., 2021). We provide the pseudocode of this setup in Appendix A. We also note that in Appendix C, we present results of performing ope using fqe instead lspe, and find that krope still reliably produces stable ope estimates.
We consider the following three classes (and total $6$) auxiliary representation learning algorithms that are typically paired with fqe for stability: I) bisimulation-metric based, II) model-based, and III) co-adaptation penalty-based. In class I, we consider 1) krope (ours), 2) deep bisimulation for control (dbc) (Zhang et al., 2021), 3) representations for ope (rope) (Pavse & Hanna, 2023a; Castro et al., 2021); in class II, we consider bellman complete representation learning (bcrl-exp-na) (Chang et al., 2022); and in class III, we consider 1) absolute dr3 regularizer (Kumar et al., 2022; Ma et al., 2024) and 2) beer regularizer (He et al., 2024). In all cases, the penultimate layer features of fqe’s action-value encoder $\phi_{\xi^{\prime}}$ are fed into lspe for ope. Note that since bcrl was not designed as an auxiliary task (Chang et al., 2022), we evaluate it as a non-auxiliary (na) task algorithm. We provide additional details on the baselines in Appendix C.